The larger of two numbers is 9 more than 3 times the smaller. If 2 times the larger is taken from 9 times the smaller, the result is 15. Find both numbers

Answer by stanbon(75887)

The larger of two numbers is 9 more than 3 times the smaller. If 2 times the larger is taken from 9 times the smaller, the result is 15. Find both numbers
------
Equations:
L = 3*S + 9
9S - 2L = 15
-----
Substitute for "L" and solve for "S"::
9S - 2(3S+9) = 15
9S - 6S - 18 = 15
3S = 33
----
Smaller = 11
---
Larger = 3*S+9 = 33+9 = 42
-------------------------------
Cheers,
Stan H.
--------